# LiDAR-based End-to-end Temporal Perception for Vehicle-Infrastructure Cooperation

### v1

@@ Abstract
Temporal perception, the ability to detect and track objects over time, is critical in autonomous driving for maintaining a comprehensive understanding of dynamic environments. However, this task is hindered by significant challenges, including incomplete perception caused by occluded objects and observational blind spots, which are common in single-vehicle perception systems. To address these issues, we introduce LET-VIC, a LiDAR-based End-to-End Tracking framework for Vehicle-Infrastructure Cooperation (VIC). LET-VIC leverages Vehicle-to-Everything (V2X) communication to enhance temporal perception by fusing spatial and temporal data from both vehicle and infrastructure sensors. First, it spatially integrates Bird’s Eye View (BEV) features from vehicle-side and infrastructure-side LiDAR data, creating a comprehensive view that mitigates occlusions and compensates for blind spots. Second, LET-VIC incorporates temporal context across frames, allowing the model to leverage historical data for enhanced tracking stability and accuracy. To further improve robustness, LET-VIC includes a Calibration Error Compensation (CEC) module to address sensor misalignments and ensure precise feature alignment. Experiments on the V2X-Seq-SPD dataset demonstrate that LET-VIC significantly outperforms baseline models, achieving at least a 13.7% improvement in mAP and a 13.1% improvement in AMOTA without considering communication delays. This work offers a practical solution and a new research direction for advancing temporal perception in autonomous driving through vehicle-infrastructure cooperation.
@@ I Introduction
In this study, we propose LET-VIC, the first LiDAR-based End-to-end Tracking framework for VIC, which designed to address major perception challenges in autonomous driving, including limited perception range and occlusion. LET-VIC fully utilizes LiDAR point clouds from both vehicle-side and infrastructure-side sensors, incorporating temporal data to greatly enhance object detection and tracking in complex traffic environments. The framework features an advanced feature fusion strategy that uses a VIC cross-attention mechanism to seamlessly integrate Bird’s Eye View (BEV) features from the infrastructure-side into the vehicle-side BEV representation. Furthermore, We employ a Calibration Error Compensation (CEC) module to compensate for calibration errors between vehicle-side and infrastructure-side sensors, delivering precise and robust 3D object perception. Rigorous testing on the V2X-Seq-SPD dataset [9], which captures challenging real-world driving scenarios, demonstrates that LET-VIC consistently outperforms existing cooperative tracking methods.
@@ II-A Cooperative Perception
With advancements in Vehicle-to-Everything (V2X) communication technologies [10, 11, 12, 13], cooperative perception in autonomous driving enhances perception capabilities through data sharing and fusion across multiple agents. This approach addresses challenges such as occlusion, long-distance perception, and sensor failures that single-vehicle perception may struggle with. Cooperative perception research primarily focuses on two types: Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) cooperation.
In V2V cooperation, several methodologies have been explored. Early Fusion techniques, as seen in Cooper [14] and DiscoNet [15], integrate raw sensor data from multiple vehicles, improving detection accuracy and extending perception range. Feature Fusion approaches, such as OPV2V [16], F-Cooper [17], V2VNet [18], and COOPERNAUT [19], combine intermediate-level features from multiple vehicles for enhanced perception. Notably, How2comm [20] focuses on transmitting BEV features extracted from point clouds to enhance cooperative 3D object detection.
In contrast, V2I cooperation has been explored through models such as V2X-ViT [21], FFNet [22], CTCE [23] and V2X-Graph [24], which adopt a middle-fusion approach to integrate data from both vehicles and infrastructure. These models account for challenges like latency and calibration errors, primarily focusing on object detection tasks. While V2V cooperation benefits from richer shared data among vehicles, V2I is especially effective in scenarios where infrastructure sensors complement vehicle perception, providing broader environmental awareness.
@@ II-B End-to-End Temporal Perception
The paradigm of End-to-End Temporal Perception has redefined sequential object perception by shifting from traditional tracking-by-detection to a more integrated approach. This approach combines object detection and tracking within a unified model. Models such as MotionTrack [4], TrackFormer [5], MOTR [6], CenterTrack [7], and CenterPoint [8] exemplify this shift, with each model contributing to multi-object temporal perception through different methods, such as sensor fusion, transformer-based architectures, and point tracking.
MotionTrack [4] introduces a transformer-based end-to-end temporal perception algorithm utilizing LiDAR-camera fusion, setting a new benchmark for tracking multiple object classes with multimodal sensor inputs. TrackFormer [5] and MOTR [6] reformulate temporal perception as a set prediction problem, leveraging transformer architectures to maintain track identity and spatiotemporal trajectories without requiring separate motion or appearance modeling. CenterTrack [7] simplifies the process by performing detection and tracking simultaneously, significantly improving temporal perception efficiency and accuracy. CenterPoint [8], a LiDAR-based 3D object detection and tracking framework, introduces an anchor-free approach that directly predicts object center points and regresses their attributes.
While end-to-end temporal perception offers substantial improvements in efficiency and accuracy, these existing methods generally lack LiDAR-based multi-view information fusion, which can limit their perception capabilities in complex environments. To address this gap, we propose a novel LiDAR-based end-to-end temporal perception framework for VIC that leverages multi-view fusion between vehicle-side and infrastructure-side sensors. This framework enhances perception by integrating complementary spatial and temporal information, enabling more accurate and robust temporal perception, demonstrating a marked advancement in the domain of cooperative temporal perception.
@@ II-C BEV-based Perception
BEV-based perception has emerged as a powerful method for enhancing object detection and tracking in autonomous driving. LSS [25] pioneered the use of depth estimation to project image features into BEV space, effectively addressing the challenge of fusing multi-view image data. BEVDet4d [26] further optimized this process, while BEVDepth [27] and BEVStereo [28] improved the precision of image-to-BEV feature projection. Methods like Simple-BEV [29], Fast-BEV [30], and MatrixVT [31] made BEV feature generation more efficient, while BEVFormer [32] introduced spatial cross-attention to fuse features from different camera views in BEV space. These advancements laid the foundation for more sophisticated end-to-end autonomous driving systems [33, 34]. Additionally, recent works such as Is-Fusion [35], BEVFusion [36], and FusionFormer [37] have focused on multi-modal data integration, combining image and LiDAR data for more accurate perception. BEV-based perception offers significant advantages for autonomous driving by enabling accurate, unified spatial representations from multi-view and multimodal data. We adopt BEV-based perception as it allows our system to integrate multi-sensor inputs seamlessly and accurately, supporting more reliable and robust performance in dynamic driving environments.
Figure 3: Architecture of LET-VIC. The diagram illustrates the steps involved in the LET-VIC framework. (a) and (b) Both the infrastructure-side and vehicle-side employ PointPillars to extract LiDAR point cloud features. Then the infrastructure-side features are transmitted to the vehicle side through V2X communication and integrated into the VIC cross-attention module along with the vehicle-side features. (c) The BEV encoder of LET-VIC is inspired by BEVFormer. (d) The decoder layers of LET-VIC employ TrackFormer. (e) The VIC Cross-Attention module fuses the Bird’s Eye View (BEV) features from both the infrastructure and vehicle sides, compensating for calibration errors.
@@ III Method
This section describes the LET-VIC method utilized in our proposed LiDAR-based end-to-end tracking framework for VIC. At the core of LET-VIC is the VIC Cross-Attention mechanism, which dynamically integrates LiDAR features from both vehicle-side and infrastructure-side perspectives, with a focus on each point within the ego vehicle’s surrounding area. As illustrated in Fig. 3, LET-VIC leverages PointPillars [38] for feature extraction from both vehicle and infrastructure LiDAR data. The features are then fused using the VIC Cross-Attention module to enhance feature integration on the vehicle side, followed by detection and tracking using the MOTR approach [6]. Additionally, the framework addresses the significant impact of sensor calibration errors on detection and tracking accuracy through an online sensor calibration compensation module. We begin by defining the VIC temporal perception problem and providing relevant background information. Following this, we present our approach to generating BEV features and detail the methodology for effectively fusing multi-view data. We then address the significant challenge of sensor calibration errors and propose a compensation mechanism to mitigate their impact. Subsequently, we present our end-to-end temporal perception strategy. Finally, we discuss the evaluation metrics employed to assess the overall effectiveness of LET-VIC.
@@ III-B BEV Feature Generation
Fig. 4 illustrates how the PointPillars framework [38] is employed on both vehicle and infrastructure sides to process LiDAR point clouds. The raw point cloud data is first organized into pillars—grids that aggregate spatial information—enabling efficient feature extraction for subsequent processing.
@@ III-C BEV Feature Transmission
In V2X communication, three primary data types are transmitted: raw data, instance-level perception data, and feature-level intermediate data. Each type incurs different transmission costs and offers varying levels of information retention.
@@ III-D BEV Feature Fusion
The ability to adaptively focus on critical features enables the deformable attention mechanism to excel in complex, multi-view tracking tasks where spatial alignment may vary due to sensor calibration errors. Fig. 5 illustrates the VIC Cross-Attention mechanism, showcasing its ability to dynamically integrate data across different sensors. The process of VIC Cross-Attention (VICCrossAttn) can be formulated as follows:
@@ III-E Calibration Error Compensation
Calibration errors in sensors, arising from precision variations, environmental factors, and calibration method uncertainties, pose significant challenges to perception accuracy in autonomous driving systems. Compensating for sensor calibration errors is crucial for LET-VIC as it ensures accurate data fusion and reliable cooperative perception performance. Therefore, we implemented CEC module to optimize spatial error correction during coordinate transformation, delivering precise and robust temporal perception.
Initially, the positions of vehicle-side or infrastructure-side features corresponding to each point in the BEV features are determined through coordinate system transformation. Subsequently, CEC employs a set of learnable parameters that adapt during training to learn adaptive offsets, which include ${\Delta x}^{inf}$, ${\Delta y}^{inf}$, ${\Delta x}^{veh}$, and ${\Delta y}^{veh}$. Finally, the positions of vehicle-side or roadside features corresponding to each BEV feature are adjusted by adding this set of offsets, forming compensated reference points as shown in Eqs. 4 and  5. Then, we normalize the original reference point coordinates according to the specified point cloud range, as detailed in Eqs. 6 and  7.
@@
Where $x_{i}^{veh}$, $y_{i}^{veh}$ represents the $i$-th original coordinates of the reference point on the vehicle-side and $x_{i}^{inf}$, $y_{i}^{inf}$ represents the $i$-th original coordinates of the reference point on the infrastructure-side. $R_{bev2inf}$ represents the rotation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of infrastructure-side point cloud BEV features, and $T_{bev2inf}$ represents the translation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of infrastructure-side point cloud BEV features. ${\Delta x}_{i}^{inf}$ and ${\Delta y}_{i}^{inf}$ represent the calibration errors corresponding to the infrastructure side at $i$-th reference point; ${Ref}_{i}^{veh}$ represents the $i$-th reference points on the pseudo-image of vehicle side point cloud BEV features, $R_{bev2veh}$ represents the rotation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of vehicle side point cloud BEV features, and $T_{bev2veh}$ represents the translation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of vehicle side point cloud BEV features; ${\Delta x}_{i}^{veh}$ and ${\Delta y}_{i}^{veh}$ represent the calibration errors corresponding to the vehicle side at $i$-th reference point. $x_{min}^{v}eh$, $x_{max}^{v}eh$ are the minimum and maximum values of the vehicle side point cloud range in the x-direction, $y_{min}^{v}eh$, $y_{max}^{v}eh$ are the minimum and maximum values of the vehicle side point cloud range in the y-direction. $x_{min}^{i}nf$, $x_{max}^{i}nf$ are the minimum and maximum values of the infrastructure side point cloud range in the x-direction, $y_{min}^{i}nf$, $y_{max}^{i}nf$ are the minimum and maximum values of the infrastructure side point cloud range in the y-direction. ${P}_{i}^{veh}$, ${P}_{i}^{inf}$ are the input parameters defined in Eq. III-D.

### v2

@@ Abstract
Temporal perception, defined as the capability to detect and track objects across temporal sequences, serves as a fundamental component in autonomous driving systems. While single-vehicle perception systems encounter limitations, stemming from incomplete perception due to object occlusion and inherent blind spots, cooperative perception systems present their own challenges in terms of sensor calibration precision and positioning accuracy. To address these issues, we introduce LET-VIC, a LiDAR-based End-to-End Tracking framework for Vehicle-Infrastructure Cooperation (VIC). First, we employ Temporal Self-Attention and VIC Cross-Attention modules to effectively integrate temporal and spatial information from both vehicle and infrastructure perspectives. Then, we develop a novel Calibration Error Compensation (CEC) module to mitigate sensor misalignment issues and facilitate accurate feature alignment. Experiments on the V2X-Seq-SPD dataset demonstrate that LET-VIC significantly outperforms baseline models. Compared to LET-V, LET-VIC achieves +15.0% improvement in mAP and a +17.3% improvement in AMOTA. Furthermore, LET-VIC surpasses representative Tracking by Detection models, including V2VNet, FFNet, and PointPillars, with at least a +13.7% improvement in mAP and a +13.1% improvement in AMOTA without considering communication delays, showcasing its robust detection and tracking performance. The experiments demonstrate that the integration of multi-view perspectives, temporal sequences, or CEC in end-to-end training significantly improves both detection and tracking performance. All code will be open-sourced.
@@
This work has been accepted for publication in IEEE Internet of Things Journal.
DOI: 10.1109/JIOT.2025.3552526
@@ I Introduction
In this study, we propose LET-VIC, the first LiDAR-based End-to-end Tracking framework for VIC, which designed to address major perception challenges in autonomous driving, including limited perception range and occlusion. The core innovation lies in simultaneously enhancing perception through two complementary dimensions: spatially fusing infrastructure perspectives with vehicle sensors to expand coverage, and temporally integrating historical observations to resolve occlusion ambiguity, both of which are unified within one efficient framework. LET-VIC fully utilizes LiDAR point clouds from both vehicle-side and infrastructure-side sensors, incorporating temporal data to greatly enhance object detection and tracking in complex traffic environments. The framework features an advanced feature fusion strategy that uses a VIC cross-attention mechanism to seamlessly integrate Bird’s Eye View (BEV) features from the infrastructure-side into the vehicle-side BEV representation. This dual-fusion approach, which combines cross-sensor spatial fusion with temporal evolution modeling, enables the system to overcome both physical obstructions and single-frame sensing limitations. Furthermore, We employ a Calibration Error Compensation (CEC) module to compensate for calibration errors between vehicle-side and infrastructure-side sensors, delivering precise and robust 3D object perception. Compared to existing approaches, LET-VIC is novel in its ability to unify spatial and temporal data fusion in an end-to-end manner. This ensures not only better detection and tracking performance but also improved robustness to environmental complexity and sensor misalignments, making it a critical step forward for autonomous driving applications.
@@ II-A Cooperative Perception
Cooperative perception in autonomous driving enhances perception capabilities to address challenges such as occlusion, long-distance perception, and sensor failures that single-vehicle perception may struggle with. Recent cooperative perception research primarily focuses on two types: Vehicle-to-Vehicle (V2V) and Vehicle-to-Infrastructure (V2I) cooperation [10, 11, 12, 13].
In V2V cooperation, different fusion approaches have been explored. Early Fusion techniques, as seen in Cooper [14] and DiscoNet [15], integrate raw sensor data from multiple vehicles, improving detection accuracy and extending perception range. Feature Fusion approaches, such as OPV2V [16], F-Cooper [17], V2VNet [18], and COOPERNAUT [19], combine intermediate-level features from multiple vehicles for enhanced perception. Notably, How2comm [20] focuses on transmitting BEV features extracted from point clouds to enhance cooperative 3D object detection.
On the other hand, V2I cooperation has been explored through models such as V2X-ViT [21], FFNet [22], CTCE [23], V2X-Graph [24] and UniV2Xyu2024end, which adopt a feature fusion approach to integrate data from both vehicles and infrastructure to provide broader environmental awareness.
@@ II-B End-to-End Temporal Perception
The paradigm of End-to-End Temporal Perception has redefined sequential object perception by shifting from traditional tracking-by-detection to a unified model which has been proven to simultaneously enhance both detection and tracking accuracy [6, 25, 26, 27, 28, 29]. MotionTrack [4] introduces a transformer-based end-to-end temporal perception algorithm utilizing LiDAR-camera fusion, setting a new benchmark for tracking multiple object classes with multimodal sensor inputs. TrackFormer [5] and MOTR [6] reformulate temporal perception as a set prediction problem, leveraging transformer architectures to maintain track identity and spatiotemporal trajectories without requiring separate motion or appearance modeling. CenterTrack [7] simplifies the process by performing detection and tracking simultaneously, significantly improving temporal perception efficiency and accuracy. CenterPoint [8], a LiDAR-based 3D object detection and tracking framework, introduces an anchor-free approach that directly predicts object center points and regresses their attributes.
However, the end-to-end framework has not been thoroughly explored in the LiDAR-based VIC field. HYDRO-3D [30] improve detection accuracy by leveraging temporal information and collaborative perception, yet they lack a comprehensive evaluation of tracking performance. MOT-CUP [31] proposes a joint perception post-processing framework validated under ideal simulation V2X-Sim dataset [32]. However, the effectiveness of real-time collaborative perception and tracking is significantly impacted by transmission delays, data timestamp alignment, and the precision of coordinate transformation matrices, all of which are critical and non-negligible factors. These issues substantially increase the complexity of hyperparameter tuning for such post-processing algorithms. In contrast, end-to-end training addresses these challenges by employing sophisticated network designs, eliminating the need for manual hyperparameter configuration while achieving greater robustness. [33] and [34] introduce novel 3D multi-object cooperative tracking algorithms based on a post-fusion framework, evaluated on the real-world V2V4Real dataset [35]. However, these approaches solely focus on optimizing the tracking component and fail to improve detection accuracy as same as MOT-CUP [31].
To bridge these gaps, we introduce a novel LiDAR-based end-to-end temporal perception framework for VIC, which leverages multi-view fusion between vehicle-side and infrastructure-side sensors. By integrating complementary spatial and temporal information, this framework significantly enhances perception accuracy and robustness, marking a significant advance in cooperative temporal perception.
@@ II-C BEV-based Perception
BEV-based perception has emerged as a powerful method for enhancing object detection and tracking in autonomous driving by enabling accurate, unified spatial representations from multi-view and multimodal data. LSS [36] pioneered the use of depth estimation to project image features into BEV space, effectively addressing the challenge of fusing multi-view image data. BEVDet4d [37] further optimized this process, while BEVDepth [38] and BEVStereo [39] improved the precision of image-to-BEV feature projection. Unlike the BEVDet series of methods, BEVFormer [40] introduced spatial cross-attention to fuse features from different camera views in BEV space. Methods like Simple-BEV [41], Fast-BEV [42], and MatrixVT [43] have made BEV feature generation more efficient. Additionally, recent works such as Is-Fusion [44], BEVFusion [45], and FusionFormer [46] have focused on multi-modal data integration, combining image and LiDAR data for more accurate perception. These advancements laid the foundation for more sophisticated end-to-end autonomous driving systems [47, 48, 49, 50, 51, 52]. We adopt BEV-based perception as it allows our system to integrate multi-sensor inputs seamlessly and accurately, supporting more reliable and robust performance in dynamic driving environments.
Figure 3: Architecture of LET-VIC. The diagram illustrates the steps involved in the LET-VIC framework. (a) and (b) Both the infrastructure-side and vehicle-side employ PointPillars [53] to extract LiDAR point cloud features. Then the infrastructure-side features are transmitted to the vehicle side through V2X communication and integrated into the VIC cross-attention module along with the vehicle-side features. (c) The BEV encoder of LET-VIC is inspired by BEVFormer. (d) The decoder layers of LET-VIC employ TrackFormer. (e) The VIC Cross-Attention module fuses the Bird’s Eye View (BEV) features from both the infrastructure and vehicle sides, compensating for calibration errors.
@@ III Method
This section describes the LET-VIC method utilized in our proposed LiDAR-based end-to-end tracking framework for VIC. LET-VIC’s key novelty lies in dynamically combining LiDAR data from both vehicle and infrastructure perspectives through a novel VIC Cross-Attention mechanism. This design enables more accurate object detection and tracking by addressing challenges such as sensor calibration errors and limited individual sensor coverage. In essence, LET-VIC enhances environmental awareness, enabling safer and more efficient autonomous driving. The end-to-end design further simplifies the traditional tracking pipeline, combining object detection and tracking into a seamless, jointly optimized framework, improving efficiency and accuracy. As illustrated in Fig. 3, LET-VIC leverages PointPillars [53] for feature extraction from both vehicle and infrastructure LiDAR data. The features are then fused using the VIC Cross-Attention module to enhance feature integration on the vehicle side, followed by detection and tracking using the MOTR approach [6]. Additionally, the framework addresses the significant impact of sensor calibration errors on detection and tracking accuracy through an online sensor calibration compensation module. We begin by defining the VIC temporal perception problem and providing relevant background information. Following this, we present our approach to generating BEV features and detail the methodology for effectively fusing multi-view data. We then address the significant challenge of sensor calibration errors and propose a compensation mechanism to mitigate their impact. Subsequently, we present our end-to-end temporal perception strategy. Finally, we discuss the evaluation metrics employed to assess the overall effectiveness of LET-VIC.
@@ III-B BEV Feature Generation
Figure 4 provides a detailed illustration of the Pointcloud Backbone used in both the vehicle-side and infrastructure-side components of the LET-VIC framework, corresponding to parts (a) and (b) in Fig. 3. The diagram shows how the PointPillars framework [53] is employed to process LiDAR point clouds on both sides. The raw point cloud data is first organized into pillars—grids that aggregate spatial information—enabling efficient feature extraction for subsequent processing.
@@ III-C BEV Feature Transmission
The block diagram between parts (a) and (b) of Fig. 3 illustrates the transmission of BEV features from the infrastructure side to the vehicle side. In V2X communication, three primary data types are transmitted: raw data, instance-level perception data, and feature-level intermediate data. Each type incurs different transmission costs and offers varying levels of information retention.
@@ III-D BEV Feature Fusion
The ability to adaptively focus on critical features enables the deformable attention mechanism to excel in complex, multi-view tracking tasks where spatial alignment may vary due to sensor calibration errors. Figure 5 corresponds to part (e) in Fig. 3 and illustrates the VIC Cross-Attention mechanism, showcasing its ability to dynamically integrate data across different sensors. The process of VIC Cross-Attention (VICCrossAttn) can be formulated as follows:
@@ III-E Calibration Error Compensation
Calibration errors in sensors, arising from precision variations, environmental factors, and calibration method uncertainties, pose significant challenges to perception accuracy in autonomous driving systems. Compensating for sensor calibration errors is crucial for LET-VIC as it ensures accurate data fusion and reliable cooperative perception performance. Therefore, we implemented CEC module to optimize spatial error correction during coordinate transformation, delivering precise and robust temporal perception. The CEC module is integrated into the VIC Cross-Attention mechanism, which corresponds to part (e) in Fig. 3.
Initially, the positions of vehicle-side or infrastructure-side features corresponding to each point in the BEV features are determined through coordinate system transformation. Subsequently, CEC employs a set of learnable parameters that adapt during training to learn adaptive offsets, which include ${\Delta x}^{inf}$, ${\Delta y}^{inf}$, ${\Delta x}^{veh}$, and ${\Delta y}^{veh}$. Finally, the positions of vehicle-side or roadside features corresponding to each BEV feature are adjusted by adding this set of offsets, forming compensated reference points as shown in (4) and (5). Then, we normalize the original reference point coordinates according to the specified point cloud range, as detailed in (6) and (7).
@@
Where $x_{i}^{veh}$, $y_{i}^{veh}$ represents the $i$-th original coordinates of the reference point on the vehicle-side and $x_{i}^{inf}$, $y_{i}^{inf}$ represents the $i$-th original coordinates of the reference point on the infrastructure-side. $R_{bev2inf}$ represents the rotation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of infrastructure-side point cloud BEV features, and $T_{bev2inf}$ represents the translation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of infrastructure-side point cloud BEV features. ${\Delta x}_{i}^{inf}$ and ${\Delta y}_{i}^{inf}$ represent the calibration errors corresponding to the infrastructure side at $i$-th reference point; ${Ref}_{i}^{veh}$ represents the $i$-th reference points on the pseudo-image of vehicle side point cloud BEV features, $R_{bev2veh}$ represents the rotation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of vehicle side point cloud BEV features, and $T_{bev2veh}$ represents the translation matrix from the vehicle side LiDAR coordinate system to the pseudo-image coordinate system of vehicle side point cloud BEV features; ${\Delta x}_{i}^{veh}$ and ${\Delta y}_{i}^{veh}$ represent the calibration errors corresponding to the vehicle side at $i$-th reference point. $x_{min}^{v}eh$, $x_{max}^{v}eh$ are the minimum and maximum values of the vehicle side point cloud range in the x-direction, $y_{min}^{v}eh$, $y_{max}^{v}eh$ are the minimum and maximum values of the vehicle side point cloud range in the y-direction. $x_{min}^{i}nf$, $x_{max}^{i}nf$ are the minimum and maximum values of the infrastructure side point cloud range in the x-direction, $y_{min}^{i}nf$, $y_{max}^{i}nf$ are the minimum and maximum values of the infrastructure side point cloud range in the y-direction. ${P}_{i}^{veh}$, ${P}_{i}^{inf}$ are the input parameters defined in (III-D).